17 Feb 2019 UT
Plan: Start with LUCI1/LUCI2
Dither test for Bino LUCI and ESM LUCI2. See if the detxy +y dithers move up/down along the slit, or is there an offset?
Collect Pointing Model data with collimation & then LUCI images
Open-loop tracking test (deferred)
Weather: Although it has been clearing up, fog remains over the mountain. Delayed opening due to fog.
~01:00 UT: L2 MOS error during "initialize all". Strain gauges. Jenny fixed it.
01:50 UT: Set up to take L2 Ks flats for Pinna AO program. N30 N30_FS.
01:53 UT: Another L2 MOS error. This one occurred while moving from "no mask" to the N30_FieldStop. The error is in the sequence MOS - storage to turnout.
02:16 UT: Jenny fixed it; I initialized MOS. Grabber failing to trigger positive limit switch - not grabbing the mask. Failed again. Recovery will take ~1 hr.
~02:46 UT: Tel @ zenith and lights off. Taking dual imaging focus sequences with MODS1 and MODS2. MODS2b sieveSnaps look a bit elongated. mods??.20170217.0002---0012.fits
Note added 18-Feb-2019: I changed the best-focus for mods2b dual imaging from 1225 -> 1255, as this was the one that was most obviously out of focus on the sieve Snap images. The other values appear to be off as well, but I'll check these and the other sequences taken last night before making any other changes.
|| best-foc img #
|| best-focus [mic]
|| actual value
03:19 UT: Jenny called and said that the L2 MOS problem was resolved. MOS was exercised a lot. It is at N30_FS.
03:50 UT: Having aligned the FS, I started the N30 Ks flats. mask=N30_FS, filter=Ks, camera=N30, mirror.
|| lamp status
|| DIT x NDIT x NEXPO
|| Halo1 OFF
|| 2.8s x 2 x 5
|| Halo1 ON
|| same as above
04:50 UT: Continue MODS focus sequences.
- Blue direct imaging.
- Red direct imaging.
- Dual grating
- 17 Feb 2019